OpenStudy (anonymous):

Consider the solid S described below. The base of S is a circular disk with radius 2r. Parallel cross-sections perpendicular to the base are squares. Find the volume V of this solid.

OpenStudy (anonymous):

since all the cross sections are squares, the height of the solid, known as the "cylinder" is equal in length to the circumference of the base. Thus the volume is= pi(2r)^2*2r=pi*4r^2*2r=pi*8r^3

OpenStudy (anonymous):

for some reason this answer doesnt work...

OpenStudy (anonymous):

well, i'm probably wrong. I think this has to do with calculus and finding integrals. Is this problem for calculus?

OpenStudy (anonymous):

yes, this is calculus 2

OpenStudy (anonymous):

try this, it can help http://answers.yahoo.com/question/index?qid=20101007174749AAZyxDi